# On arithmetic terms expressing the prime-counting function and the n-th prime

### v1

@@ 3 Singlefold definitions for binomial coefficients
We denote this version of the singlefold relation $z=\binom{x}{y}$ as
@@
This version contains $12$ quantified variables. Reducing the number of quantified variables can vastly decrease the magnitude of arithmetic terms constructed using the hypercube method. The singlefold relation $z=\binom{x}{y}$ is important and will be used in the definition of additional singlefolds, so it would be ideal if we could see some reduction here.
Indeed, we find a new arithmetic term for $z=\binom{x}{y}$ that allows us to reduce the number of quantified variables from $12$ to $7$. This new version will be written as
@@ 3.1 A new arithmetic term for binomial coefficients
So, the row number $7$ of the matrix contains the row number $6$ of Pascal’s Triangle, while the $64$-th element $s_{8}(63)$ is equal to $2$, as from now on the row-sides collide and the simulation ends. To illustrate:
@@
3.2 An improved singlefold definition for binomial coefficients
Since all exponentiations in the arithmetic terms from Theorem 3.1 and Theorem 3.2 are performed using powers of $2$, we no longer require the singlefold relation $E_{\exp}(x,y,[4],z)=0$ in our construction of the singlefold for $z=\binom{x}{y}$. This results in an immediate reduction of $4$ quantified variables. We get a further reduction of $1$ quantified variable, for a total reduction of $5$ quantified variables:
@@ 3.2 An improved singlefold definition for binomial coefficients
4 Singlefold definitions for factorials and Hamming weights
We will now develop our singlefold relations for the factorial and Hamming weight functions. We start with some lemmas:
@@ 4 Singlefold definitions for factorials and Hamming weights
For $n=0$ we have $1\geq 1$. For $n=1$ we have $8\geq 8$. For $n=2$, we have $8^{4}\geq 3^{4}$. Further, the functions are more and more apart, because for $n=3$, one has already $n^{2}>(n+2)\log_{8}(n+1)$, and $n^{2}$ increases faster than the right-hand side. ∎
@@
We define the singlefold exponential Diophantine relation $f=n!$ as:
@@
A further application of the binomial coefficient is the singlefold definition of the Hamming Weight of a number $y=\operatorname{HW}(x)$. We write this expression as
@@
5 The prime omega function and square roots of unity
@@
    In $\mathbb{Z}/2\mathbb{Z}$, the congruence $x^{2}\equiv 1\pmod{2}$ has exactly $1$ solution:
@@
    In $\mathbb{Z}/4\mathbb{Z}$, the congruence $x^{2}\equiv 1\pmod{4}$ has exactly $2$ solutions:
@@
    In $\mathbb{Z}/2^{k}\mathbb{Z}$ for $k\geq 3$, the congruence $x^{2}\equiv 1\pmod{2^{k}}$ has exactly $4$ solutions:
@@
    In $\mathbb{Z}/p^{k}\mathbb{Z}$ for $p$ an odd prime and $k\geq 1$, the congruence $x^{2}\equiv 1\pmod{p^{k}}$ has exactly $2$ solutions:
@@
(2 solutions in /4)×(2 solutions in /$(2\text{ solutions in }\mathbb{Z}/4\mathbb{Z})\times(2\text{ solutions in }\mathbb{Z}/p_{1}^{b_{1}}\mathbb{Z})\times\cdots\times(2\text{ solutions in }\mathbb{Z}/p_{d}^{b_{d}}\mathbb{Z})=2^{d+1}.$
@@
In this case, $n=2^{a}p_{1}^{b_{1}}\cdots p_{d}^{b_{d}}$, $a\geq 1$ and $\omega(n)=d+1$. The exponent of $2$ in $4n$ will be $a+2\geq 3$, so the number of solutions in $\mathbb{Z}/(4n)\mathbb{Z}$ will be:
(4 solutions in /$(4\text{ solutions in }\mathbb{Z}/2^{a+2}\mathbb{Z})\times(2\text{ solutions in }\mathbb{Z}/p_{1}^{b_{1}}\mathbb{Z})\times\cdots\times(2\text{ solutions in }\mathbb{Z}/p_{d}^{b_{d}}\mathbb{Z})=2^{d+2}.$
@@
We see that the number of distinct prime divisors of a positive integer $n$ can be determined by a fixed-length elementary closed form expression. Most interesting is that this expression does not reveal any information about the specific divisors of $n$. Recall that $M(n)$ Equation 17 is an fixed-length elementary closed form expression and so are $\nu_{2}(n)$ and $\operatorname{HW}(n)$ as shown in § 2.1.
@@ 6 The prime-counting function
The proof is trivial, though we note that the factorial function $n!=n(n-1)!$ defines $0!=1$. ∎
@@ 7 An exponential Diophantine equation for the n-th prime
Of considerable importance will be the arithmetic term $M(n)$ Equation 17, which we used in previous sections to develop our arithmetic terms for $\mathcal{N}(n)$, $\omega(n)$, and $\pi(n)$. We will require the singlefold relation $m=M(4n)$, which will be written as:
@@
The first equation, $E_{!}(x_{1},[13],x_{2})=0$, defines $x_{2}=x_{1}!$ singlefold.
The second equation, $E_{M4}(x_{2},[0],x_{3})=0$, defines $x_{3}=M(4x_{2})=M(4(x_{1}!))$ singlefold.
The third equation, $E_{\operatorname{HW}}(x_{3},[12],x_{4})=0$, defines $x_{4}=\operatorname{HW}(x_{3})=\operatorname{HW}(M(4(x_{1}!)))$ singlefold.
@@
Therefore, for all $n\in\mathbb{Z}^{+}$, the number of solutions to the system is generated by the set $\{\vec{x}\in\mathbb{N}^{29}\}$. ∎
@@ 7.1 Building the equation
During the construction of the equation $F(n,\vec{x})=0$, we will relabel some of the $x_{i}$ variables in the original system to make it easier to distinguish which variables arise from certain singlefold relations.
@@
The final equation $F(n,\vec{x},[25])=0$ is a sum of squares of the equations defined in Lemma 7.3, which are all singlefold relations. Furthermore, the values of all singlefold relations in the set of equations are uniquely determined by the variable $x_{1}$. Hence, by Lemma 2.1, the solution set of the system coincides with the solution set $\{(\vec{x},\vec{y})=(x_{1},x_{2},\ldots,x_{7},y_{1},y_{2},\ldots,y_{25})\in\mathbb{N}^{32}\}:F(n,\vec{x},[25])=0$. ∎
@@ 8 Reducing the number of monomials
The arithmetic term $Q(n)$ arises as the sum of a single $\mathcal{C}_{32}(m_{i}(n,\vec{x}),t(n),u(n))$ term and $10099$ distinct $\mathcal{A}_{32}(m_{i}(n,\vec{x}),t(n),u(n))$ terms. Every $\mathcal{A}_{32}(m_{i}(n,\vec{x}),t(n),u(n))$ term is the product of $32$ $G_{r}(b,t(n))$ terms, where $r\in\{0,\ldots,31\}$ and $b\in\mathbb{N}$ is determined by $n$ and $\vec{x}$.
Unfortunately, this arithmetic term for $p(n)$ is far too lengthy: Merely writing its $\mathcal{C}_{k}(\cdot)$ term and many $\mathcal{A}_{k}(\cdot)$ terms would occupy roughly $50$ pages in this paper. Consequently, establishing the bounds for the arithmetic terms $t(n)$, $u(n)$ would be an extremely tedious exercise and also cumbersome to verify. We opt instead to find an arithmetic term for $p(n)$ that can be written down completely on a few pages as a sum of its $\mathcal{C}_{k}(\cdot)$ term and $\mathcal{A}_{k}(\cdot)$ terms. This necessitates reducing the number of monomials in our exponential Diophantine equation $F(n,\vec{x},[25])=0$ from $10102$ down to $500$ or so. Such term will allow our results to be more easily analyzed and shared.
It is possible to significantly reduce the number of monomials by introducing new variables and breaking the large square relations down into smaller squares. However, each additional variable introduced increases the magnitude of the resulting arithmetic term, creating a trade-off. While the final $p(n)$ term we present in § 10 requires fewer $\mathcal{C}_{k}(\cdot)$ and $\mathcal{A}_{k}(\cdot)$ terms to write down, the $32$ variable version is in actuality much smaller from a computational perspective.
@@
In this equation, the vast majority of monomials arise from the expansion of the square term $E_{M4}$, which produces a staggering $9920$ of the $10102$ total monomials. Our goal now is to reduce the number of monomials in the expansion of this square by splitting it up into a sum of smaller squares.
@@
This equation contains $503$ monomials in $42$ quantified variables. For its full monomial expansion, please consult Appendix appendix A.
@@
9 Singlefold bounds
To write down an explicit arithmetic term for $p(n)$, we must find suitable arithmetic terms $t(n),u(n)$ that allow us to apply the hypercube method to count the number of solutions to our exponential Diophantine equation $\hat{F}(n,\vec{x})=0$. The first step in this process is to establish the bounds for all quantified variables in our singlefold relations.
@@ 9 Singlefold bounds
$E_{\operatorname{HW}}(x_{1},[12],x_{2})=0\iff\exists\vec{y}\in\mathbb{N}^{12}:E_{\binom{\#}{\#}}(2x_{1},x_{1},[7],y_{1})+E_{\nu}(y_{1},[4],x_{2})=0,$
@@ 10 The $n$-th prime function
$<\ell\cdot 2^{2^{9t(n)+8}}=503\cdot 2^{2^{9t(n)+8}}<2^{9}\cdot 2^{2^{9t(n)+8}}=2^{2^{9t(n)+8}+9}.$

### v2

@@ 3 Single-fold definitions for binomial coefficients
We denote this version of the single-fold relation $z=\binom{x}{y}$ as
@@
This version contains 121212 quantified variables. Reducing the number of quantified variables can vastly decrease the magnitude of arithmetic terms constructed using the hypercube method. The single-fold relation $z=\binom{x}{y}$ is important and will be used in the definition of additional single-folds, so it would be ideal if we could see some reduction here.
Indeed, we find a new arithmetic term for $z=\binom{x}{y}$ that allows us to reduce the number of quantified variables from 121212 to 777. This new version will be written as
@@ 3.1 A new arithmetic term for binomial coefficients
So, the row number 777 of the matrix contains the row number 666 of Pascal’s Triangle, while the 646464-th element $s_{8}(63)$ is equal to 222, as from now on the row-sides collide and the simulation ends. To illustrate:
@@
3.2 An improved single-fold definition for binomial coefficients
Since all exponentiations in the arithmetic terms from Theorem 3.1 and Theorem 3.2 are performed using powers of 222, we no longer require the single-fold relation $E_{\exp}(x,y,[4],z)=0$ in our construction of the single-fold for $z=\binom{x}{y}$. This results in an immediate reduction of 444 quantified variables. We get a further reduction of 111 quantified variable, for a total reduction of 555 quantified variables:
@@ 3.2 An improved single-fold definition for binomial coefficients
4 Single-fold definitions for factorials and Hamming weights
We will now develop our single-fold relations for the factorial and Hamming weight functions. We start with some lemmas:
@@ 4 Single-fold definitions for factorials and Hamming weights
For $n=0$ we have $\geq 1$ 1. For $n=1$ we have $\geq 8$ 8. For $n=2$, we have $8^{4}\geq 3^{4}$. Further, the functions are more and more apart, because for $n=3$, one has already $n^{2}>(n+2)\log_{8}(n+1)$, and $n^{2}$ increases faster than the right-hand side. ∎
@@
We define the single-fold exponential Diophantine relation $f=n!$ as:
@@
A further application of the binomial coefficient is the single-fold definition of the Hamming Weight of a number $y=\operatorname{HW}(x)$. We write this expression as
@@
5 The prime omega function and modular square roots of unity
@@
    In $\mathbb{Z}/2\mathbb{Z}$, the congruence $x^{2}\equiv 1\pmod{2}$ has exactly 111 solution:
@@
    In $\mathbb{Z}/4\mathbb{Z}$, the congruence $x^{2}\equiv 1\pmod{4}$ has exactly 222 solutions:
@@
    In $\mathbb{Z}/2^{k}\mathbb{Z}$ for $k\geq 3$, the congruence $x^{2}\equiv 1\pmod{2^{k}}$ has exactly 444 solutions:
@@
    In $\mathbb{Z}/p^{k}\mathbb{Z}$ for $p$ an odd prime and $k\geq 1$, the congruence $x^{2}\equiv 1\pmod{p^{k}}$ has exactly 222 solutions:
@@
$(2\text{ solutions in }\mathbb{Z}/4\mathbb{Z})\times(2\text{ solutions in }\mathbb{Z}/p_{1}^{b_{1}}\mathbb{Z})\times\cdots\times(2\text{ solutions in }\mathbb{Z}/p_{d}^{b_{d}}\mathbb{Z})=2^{d+1}.$
@@
[…]
$(4\text{ solutions in }\mathbb{Z}/2^{a+2}\mathbb{Z})\times(2\text{ solutions in }\mathbb{Z}/p_{1}^{b_{1}}\mathbb{Z})\times\cdots\times(2\text{ solutions in }\mathbb{Z}/p_{d}^{b_{d}}\mathbb{Z})=2^{d+2}.$
@@
We see that the number of distinct prime divisors of a positive integer $n$ can be determined by a fixed-length elementary closed-form expression. Most interesting is that this expression does not reveal any information about the specific divisors of $n$. Recall that $M(n)$ Equation 17 is a fixed-length elementary closed-form expression and so are $\nu_{2}(n)$ and $\operatorname{HW}(n)$ as shown in § 2.1.
@@ 6 The prime-counting function
The proof is trivial, though we note that the factorial function $n!=n(n-1)!$ defines 0!=10!=10 ! = 1. ∎
@@ 7 An exponential Diophantine equation for the n-th prime
Of considerable importance will be the arithmetic term $M(n)$ Equation 17, which we used in previous sections to develop our arithmetic terms for $\mathcal{N}(n)$, $\omega(n)$, and $\pi(n)$. We will require the single-fold relation $m=M(4n)$, which will be written as:
@@
The first equation, $E_{!}(x_{1},[13],x_{2})=0$, defines $x_{2}=x_{1}!$ single-fold.
The second equation, $E_{M4}(x_{2},[0],x_{3})=0$, defines $x_{3}=M(4x_{2})=M(4(x_{1}!))$ single-fold.
The third equation, $E_{\operatorname{HW}}(x_{3},[12],x_{4})=0$, defines $x_{4}=\operatorname{HW}(x_{3})=\operatorname{HW}(M(4(x_{1}!)))$ single-fold.
@@
Therefore, for all $n\in\mathbb{Z}^{+}$, the number of solutions to the system is generated by the set $\{(\vec{x},\vec{y})\in\mathbb{N}^{29}\}$. ∎
@@ 7.1 Building the equation
During the construction of the equation $F(n,\vec{x})=0$, we will relabel some of the $x_{i}$ variables in the original system to make it easier to distinguish which variables arise from certain single-fold relations.
@@
The final equation $F(n,\vec{x},[25])=0$ is a sum of squares of the equations defined in Lemma 7.3, which are all single-fold relations. Furthermore, the values of all single-fold relations in the set of equations are uniquely determined by the variable $x_{1}$. Hence, by Lemma 2.1, the solution set of the system coincides with the solution set $\{(\vec{x},\vec{y})=(x_{1},x_{2},\ldots,x_{7},y_{1},y_{2},\ldots,y_{25})\in\mathbb{N}^{32}\}:F(n,\vec{x},[25])=0$. ∎
@@ 8 Reducing the number of monomials
The arithmetic term $Q(n)$ arises as the sum of a single $\mathcal{C}_{32}(m_{i}(n,\vec{x}),t(n),u(n))$ term and 100991009910099 distinct $\mathcal{A}_{32}(m_{i}(n,\vec{x}),t(n),u(n))$ terms. Every $\mathcal{A}_{32}(m_{i}(n,\vec{x}),t(n),u(n))$ term is the product of 323232 $G_{r}(b,t(n))$ terms, where $r\in\{0,\ldots,31\}$ and $b\in\mathbb{N}$ is determined by $n$ and $\vec{x}$.
Unfortunately, this arithmetic term for $p(n)$ is far too lengthy: Merely writing its $\mathcal{C}_{k}(\cdot)$ term and many $\mathcal{A}_{k}(\cdot)$ terms would occupy roughly 505050 pages in this paper. Consequently, establishing the bounds for the arithmetic terms $t(n)$, $u(n)$ would be an extremely tedious exercise and also cumbersome to verify. We opt instead to find an arithmetic term for $p(n)$ that can be written down completely on a few pages as a sum of its $\mathcal{C}_{k}(\cdot)$ term and $\mathcal{A}_{k}(\cdot)$ terms. This necessitates reducing the number of monomials in our exponential Diophantine equation $F(n,\vec{x},[25])=0$ from 101021010210102 down to fewer than 500500500. Such term will allow our results to be more easily analyzed and shared.
It is possible to significantly reduce the number of monomials by introducing new variables and breaking the large square relations down into smaller squares. However, each additional variable introduced increases the magnitude of the resulting arithmetic term, creating a trade-off. While the final $p(n)$ term we present in § 10 requires fewer $\mathcal{C}_{k}(\cdot)$ and $\mathcal{A}_{k}(\cdot)$ terms to write down, the 323232 variable version is in actuality much smaller from a computational perspective.
@@
In this equation, the vast majority of monomials arise from the expansion of the square term $E_{M4}$, which produces a staggering 992099209920 of the 101021010210102 total monomials. Our goal now is to reduce the number of monomials in the expansion of this square by splitting it up into a sum of smaller squares.
@@
This equation contains 498498498 monomials in 424242 quantified variables. For its full monomial expansion, please consult Appendix appendix A.
@@
9 Single-fold bounds
To write down an explicit arithmetic term for $p(n)$, we must find suitable arithmetic terms $t(n),u(n)$ that allow us to apply the hypercube method to count the number of solutions to our exponential Diophantine equation $\hat{F}(n,\vec{x})=0$. The first step in this process is to establish the bounds for all quantified variables in our single-fold relations.
@@ 9 Single-fold bounds
$x_{2}=\operatorname{HW}(x_{1})\iff E_{\operatorname{HW}}(x_{1},[12],x_{2})=0\iff\exists\vec{y}\in\mathbb{N}^{12}:E_{\binom{\#}{\#}}(2x_{1},x_{1},[7],y_{1})+E_{\nu}(y_{1},[4],x_{2})=0,$
@@ 10 The $n$-th prime function
$<\ell\cdot 2^{2^{9t(n)+8}}=498\cdot 2^{2^{9t(n)+8}}<2^{9}\cdot 2^{2^{9t(n)+8}}=2^{2^{9t(n)+8}+9}.$